can you solve it by explaning step by step and can you draw the graph of them and explain throgh the graph

5. Solve sin x < cos x on (0,21). (A) 0, 51 (B) SA -21 4 (C) π 5π 4² 4 (D) 571 ,210 4 (E) (0,2TC)
Answer #1

5.) sin( ficos -> sina a cosa ; (0,2%) Sinn- cosa co Cos* = sin(9-) sina 112-*)< 0 2.cas ****, sin X=(**)< 0 frosinca esine-s>> sin(x -74) < sinnt ; n=0,11,12 --- Now, when neo n-7/4=0 27/4 = 7 when nil x=8/4 3. X=5T14 By sign chart convention + + 0